# Stronger, Fewer, & Superior: Harnessing Vision Foundation Models for Domain Generalized Semantic Segmentation

### v4

@@ 9 Algorithm of Proposed Rein
       Get batch data:$(x,y)$ $f_{0}=L_{\text{emb}}(x)$ for $i\leftarrow 1$ to $N$ do
             $f_{i}=L_{i}(f_{i-1})$ $T_{i}=A_{i}\times B_{i}$ $S_{i}=Softmax(\frac{f_{i}\times T_{i}^{\text{T}}}{\sqrt{c}})$ $\Delta\bar{f_{i}}=S_{i}(:,2:m)\times[T_{i}(2:m)\times W_{T_{i}}+b_{T_{i}}]$ $\Delta f_{i}=(\Delta\bar{f_{i}}+f_{i})\times W_{f_{i}}+b_{f_{i}}$ $Q_{i}=T_{i}\times W_{Q_{i}}+b_{Q_{i}}$ $f_{i}=f_{i}+\Delta f_{i}$
      $\mathbb{F}_{t}\subseteq\{f_{0},f_{1},\ldots,f_{N}\}$ Calculate $Q_{max}$ and $Q_{avg}$ by Eq. (9) $Q=Concat([Q_{max},Q_{avg},Q_{N}])\times W_{Q}+b_{Q}$ $\bar{y_{t}}=\mathcal{H}(\mathbb{F}_{t},Q)$ Optimize $\mathcal{H}$ and $\mathcal{R}$ by $\mathcal{L}oss(\bar{y},y)$

### v5

@@ 12 More details about VFMs
VPT, LoRA, and AdaptFormer. Based on extensive experimentation, we have optimized the implementation of PEFT methods for DINOv2, utilizing configurations that enhance performance. These methods include: 1) VPT: It is deep and has 150 tokens. 2) LoRA: Applied to the query and value MLP components, LoRA is configured with a rank of 8. Additionally, it incorporates a minimal dropout rate of 0.1%. 3) AdaptFormer: This method employs a bottleneck design with a width of 64, initialized using LoRA. Notably, it omits layer normalization.
@@ 13 Algorithm of Proposed Rein
       Get batch data:$(x,y)$
       $f_{0}=L_{\text{emb}}(x)$
       for $i\leftarrow 1$ to $N$ do
             $f_{i}=L_{i}(f_{i-1})$
             $T_{i}=A_{i}\times B_{i}$
             $S_{i}=Softmax(\frac{f_{i}\times T_{i}^{\text{T}}}{\sqrt{c}})$
             $\Delta\bar{f_{i}}=S_{i}(:,2:m)\times[T_{i}(2:m)\times W_{T_{i}}+b_{T_{i}}]$
             $\Delta f_{i}=(\Delta\bar{f_{i}}+f_{i})\times W_{f_{i}}+b_{f_{i}}$
             $Q_{i}=T_{i}\times W_{Q_{i}}+b_{Q_{i}}$
             $f_{i}=f_{i}+\Delta f_{i}$
      $\mathbb{F}_{t}\subseteq\{f_{0},f_{1},\ldots,f_{N}\}$
       Calculate $Q_{max}$ and $Q_{avg}$ by Eq. (9)
       $Q=Concat([Q_{max},Q_{avg},Q_{N}])\times W_{Q}+b_{Q}$
       $\bar{y_{t}}=\mathcal{H}(\mathbb{F}_{t},Q)$
       Optimize $\mathcal{H}$ and $\mathcal{R}$ by $\mathcal{L}oss(\bar{y},y)$